A vehicle uses 250,000 J of energy to accelerate from rest. if the vehicle has a velocity of 50 m/s now, what is its mass? if the vehicle took 100 m to get up to full speed what force did it exert?
1
Expert's answer
2020-11-25T10:50:56-0500

The mass can be found from conservation of energy: the work was spent to gain kinetic energy:


"W=\frac12mv^2,\\\space\\
m=\frac{2W}{v^2}=200\text{ kg}."

The necessary acceleration to increase the speed from 0 to 50 m/s is


"a=\frac{v^2}{2d}."

According to Newton's second law, the force is


"F=ma=\frac{mv^2}{2d}=2500\text{ N}."